Exercise 5.5. Consider the sample space S = {(x,y) € R2 : 22 + y2 <1}, with event space & suitably chosen, and with probabiliThe Answer for i is suppose to be 0.625, but I am having trouble getting that number
Answer #1

solution;

given that

the sample space equation

S = {(x,y) \epsilon R^2: ^{x2} +^{y2}\leq1}

the probability of P(E)=area(E)/area (S)

=area(E)/\pi

Solution - Given that B C D * To be Eoch hacoing Equal Orea = T(0) 4 T 4 P(B) = P(220.420) 1/4 Alen of s 77 14 P (B) = 14 Sim* The probability function is given by X(214) -1/2 13 2 P(x) 14 14 i €(x) En f(a) = *.* 144 +10) 174) +27944) Jolayn 200 yo OY2 -3+212 ор D T X X 1/2 X 14 24 Q 1/4 + 1 31 +27 = 0-65 €) = ( 18 14 ( % van * tos van de ++ tv +1+1} /0743 Jolyo di do 072(c + -48*74 + 1/273 74 + BxY4 + 16144) นาง 13h Sandy so Do 9 = { set ya * * il fondo - 94 6-4ed =12 Thi g3 g do dol do do 946=8X (205) +27 + 36 x (1.186)+54 X10:65) 121.268 V(x) € 02 - [ECX))? = 1.188-1065)? 0.7655